### v2

# Multi-View Projection for Unsupervised Domain Adaptation in 3D Semantic Segmentation

Abstract
3D semantic segmentation plays a pivotal role in autonomous driving and road infrastructure analysis, yet state-of-the-art 3D models are prone to severe domain shift when deployed across different datasets. In this paper, we propose a novel multi-view projection framework that excels in unsupervised domain adaptation (UDA). Our approach first aligns Lidar scans into coherent 3D scenes and renders them from multiple virtual camera poses to create large-scale synthetic 2D datasets (PC2D) in various modalities. We then use them to train an ensemble of 2D segmentation models in pointcloud view domain on each modality. During inference, the models process hundreds of views per scene; the resulting logits are back-projected to 3D with an occlusion-aware voting scheme to generate final point-wise labels. These labels can be used directly or to fine-tune a 3D segmentation model in the target domain. We evaluate our approach in both Real-to-Real and Simulation-to-Real UDA settings. We achieve state-of-the-art results in UDA on Real-to-Real. We also demonstrate how our method can be used to segment ‘rare classes’, for which target 3D annotations are not available, by only using 2D annotations for those classes and leveraging 3D annotations for other classes in a source domain.
I Introduction
3D semantic segmentation of Lidar scenes is crucial for applications such as autonomous driving (AD) and road infrastructure analysis. In particular, many applications require models to be able to generalize to unseen data, which may not be from the same domain as the training data, e.g. if the data was acquired in a different location or with a different sensor. The main attempt to mitigate this problem is Unsupervised Domain Adaptation (UDA).
Most existing methods are tailored for AD, thus making concessions on accuracy to guarantee real-time performance on single or few Lidar scans.
We propose a method that leverages aligned scenes and larger computation times to perform 3D semantic segmentation. In particular, it reaches state-of-the-art-level performance in UDA.
Our method belongs to the multi-view projection family, where 3D scenes are analyzed via multiple 2D representations. We position virtual cameras around aligned Lidar scenes to generate rendered views from various perspectives. While similar methods use camera images for training and inference [1], we create large-scale synthetic datasets by rendering both the 3D scenes and their corresponding ground truth annotations. This enables a 2D segmentation model to be trained on data closely matching the inference domain, allowing to freely choose the optimal parameters for selecting and rendering the views. The extensive use of 2D models leads to interesting domain adaptation performance. We demonstrate the effectiveness of our method for unsupervised domain adaptation in Real-to-Real tasks on the nuScenes [2] and SemanticKITTI [3] datasets, and in a Simulation-to-Real task on the SynLidar [4] dataset. We also demonstrate a potential application of our method to segment ‘rare classes’, where 3D annotations are not available for the desired target classes, but 2D annotations are available along with 3D annotations for other classes. Finally, we conduct a full ablation study to verify the importance of each component of our method. Our main contributions are the following:
A novel modular multi-view projection framework with dataset generation capabilities that can be used to explore many approaches and parameters for 3D semantic segmentation
A state-of-the-art method for unsupervised domain adaptation in 3D semantic segmentation
We demonstrate an application of our method to segment ‘rare classes’ (defined above)
II Related works
II-A 3D semantic segmentation
3D semantic segmentation is the task of assigning a semantic label to each point in a 3D pointcloud. It is generally done either on Lidar pointclouds [2, 3], or RGB-D images [5]. Methods are usually divided in 3D-based and 2D-based. 3D models directly operate on the pointcloud using 3D sparse convolutions [6], while 2D models re-use 2D image segmentation methods by projecting the pointcloud to a 2D plane [7]. Importantly, 3D models have been shown to suffer from domain shift
to a higher degree than 2D models [8]. 2D approaches generally project the pointcloud to a single 2D image with channels representing various features, e.g. height (in Bird’s Eye View), intensity, density, etc. Multi-view projection methods try to leverage 2D models even further by repeatedly using the same model on multiple views of the same 3D scene.
II-B Multi-view projection for segmentation
The goal of multi-view projection methods is to extract as much knowledge from a 2D model as possible for 3D segmentation. This is done by inferring on multiple views of the 3D scene and then merging the resulting 2D segmentations. The problem then decomposes itself into two parts:
1.
Which views to use for inference and with which model ?
2.
How to merge the 2D segmentation masks to get unique 3D labels ?
Notice that the second problem is equivalent to the 2D-3D label propagation problem [9]. It is generally solved using either a voting scheme [10, 11, 12], where either the 2D masks or the probabilities/logits of the model are accumulated on points as votes, or with more complex methods such as Neural Networks [13, 14, 15].
Our proposed method uses a simple voting scheme on accumulated logits from a single 2D model.
In this paper, we focus on the first problem. Initially, [16] proposed multi-view projection for indoor RGB-D scenes, where the views are simply the RGB-D images themselves. For outdoor Lidar scenes, the same reasoning led [1] to use aligned RGB camera images as views, for which 2D annotated data is available [17]. However, this approach is limited as it requires using additional sensors, and the performance is tied to the quality of the sensor alignment. Mitigating this, [12] proposed to infer on rendered views of 3D scenes with an out-of-domain 2D model trained on camera images. This has the advantage of not requiring 3D labels or additional sensors, but comes with several limitations:
The domain gap between camera images and views of Lidar pointclouds is substantial, which negatively impacts 2D model performance.
The multi-view aspect of the method is limited because the view generation method must align with the camera images, e.g., be from a road user’s perspective.
Because aligning Lidar scans is essential for obtaining dense 3D scenes, the method is limited to static classes, as dynamic classes often appear with a so-called ‘shadow effect’ in the aligned scene, which is out of the camera images’ domain.
We propose to solve all of these limitations by leveraging 3D annotations. This makes our method more dependent on 3D annotated data, but allows to reach significantly better performance and to tackle dynamic classes. Closest to our approach, [10] proposes to use 3D annotations to generate 2D synthetic datasets to train a 2D model for downstream multi-view projection for indoor RGB-D scenes. Our method distinguishes itself in its application to outdoor Lidar scenes, where much additional processing is required to obtain satisfying views, e.g. scan aligning and meshing, as well as in the choices for the feature rendering and view selection. We also propose an ensembling mechanism for handling the modality choice, and demonstrate the potential of such methods for unsupervised domain adaptation.
II-C Unsupervised Domain Adaptation
Unsupervised Domain Adaptation (UDA) aims to adapt a model trained on a labeled source domain to perform well on an unlabeled target domain , where no ground-truth labels are available. Unlike supervised adaptation, only the source labels and unlabeled target samples can be exploited, while the evaluation is performed on a disjoint set . This setting is particularly relevant for autonomous driving, where annotated data is costly to acquire, but raw scans from the deployment domain can be collected at scale.
We compare our approach to state-of-the-art UDA methods for 3D semantic segmentation, which mostly rely on self-training, temporal consistency, or compositional data augmentation. T-UDA [18] leverages temporal coherence by enforcing sequential pointcloud consistency and cross-sensor geometric alignment within a mean-teacher framework. Lidar-UDA [19] simulates varying scan patterns via random beam dropping and improves pseudo-label reliability through cross-frame consistency and self-training. CoSMix [20] introduces compositional mixing across domains, generating hybrid pointclouds to reduce the distribution gap between source and target. SALUDA [21] further improves self-training by enforcing domain-invariant geometric constraints on pseudo-labels. Finally, UniMix [22] extends mixing-based approaches with a unified framework that interpolates both features and labels across domains, enhancing robustness to domain shift.
Our method differs by relying entirely on a multi-view 2D projection pipeline to generate pseudo-labels for the target domain. These pseudo-labels are then used to fine-tune a 3D segmentation model. This fundamentally different approach allows to reach state-of-the-art performance in UDA, especially on large, structured classes.
III Method
III-A Base method
An overview of our method, called Seg_3D_by_PC2D , is shown in Figure 1. We choose to use [12] as a base 3D multi-view projection method, which we enhance and extend.
We provide a brief description of that method below. Here, a scene refers to the concatenation of multiple Lidar scans aligned in a common reference frame. Note that this concatenation yields denser scenes, but may introduce so called ‘shadow effects’ in the scene, due to dynamic classes moving during capture.
1.
Train a 2D model for 2D semantic segmentation on natural camera images.
2.
Render views of a 3D scenes, then process them with the model, which is out of domain.
3.
Project the 2D segmentations on the 3D pointcloud and accumulate the logits, as votes.
4.
For each point, assign the most voted label.
This method yields a 3D segmentation mask while using no 3D annotation at all. Our key innovation is to bridge the domain gap between camera images and rendered views of Lidar pointclouds by re-training the 2D model in pointcloud view domain, using 3D annotations. This allows to drastically improve the performance of the base method, and can still be applied without 3D annotations for the target classes, see the rare classes experiment IV-E2.
First, we use 3D data along with 3D annotations to generate large-scale synthetic datasets of views of 3D scenes along with 2D segmentation masks, as shown in Figure 2, which we use to train the 2D model. We call these datasets ‘PointCloud2D’ datasets (PC2D). The parameters of the view generation are chosen carefully, as they are the same that will be used for inference downstream. We optimize the number and type of virtual camera poses (VGO, III-C), as well as the visualization modalities used. We use multiple models trained on each modality, and use them together in as an ensemble (MMods, III-D). Additionally, we improve the base method’s projection by adding occlusion using depth maps (OCL, III-E). In the experiments, we validate all of these components in the ablation study IV-F.
III-B Training in pointcloud view domain using 3D annotations in 2D (PC2D)
In [12], there is a substantial gap between the domain of the generated views and the domain of the camera training images. The main extension we propose to the base method is to train the 2D model in a much closer domain using 3D annotations. To this end, we generate a large-scale synthetic dataset of 2D views of 3D scenes along with 2D segmentation masks, as shown in Figure 2.
Virtual camera poses are sampled around the 3D scene, which are then used to render 2D views of the scene. This way, we can generate an arbitrary number of (image, label) pairs, yielding a large-scale synthetic dataset. Training on such a synthetic dataset can be seen as a powerful data augmentation. The only limitation is that a limited number of annotated scenes will lead to samples with less diversity as the total number of training samples increases, which could lead to overfitting.
The use of PC2D datasets allows to choose the domain of the training images, which allows full freedom over the view generation steps during both training and inference e.g. generating bird’s eye view depth maps is possible, while views in [12], were limited to ‘car’ view rendering of Lidar intensity. The two following subsections describe how we optimize the camera pose generation and the modality choice to obtain the best performance.
III-C View Generation Optimization (VGO)
We propose to optimize the view generation by choosing relevant camera poses. In [12], the camera poses are all placed at a road user’s perspective to match training data. We keep this category and add 3 new categories of camera poses. All 4 types of camera poses are obtained based on the original Lidar sensor poses, which are already describing the trajectory of the ego vehicle along the scene. An illustration of the 4 types of camera pose is shown in Figure 2, in the VGO module, and we describe them below:
car_view: The camera is placed at a road user’s perspective, i.e. at the approximate height of a car (random yaw can be applied).
conic_view: The camera is placed on the basis of a large imaginary cone pointing downwards, with the apex at the Lidar sensor position. The camera points to the apex of the cone. This corresponds to a ‘drone’ or ‘lamp post’ view of the scene, and is also the closest to the views that human annotators would typically use to annotate a scene.
top_view: The camera is placed directly above the Lidar position, looking downwards, with variable height and roll. This corresponds to the usual ‘bird’s eye view’ of the scene.
bottom_view: The camera is placed directly below the Lidar position, looking upwards. This is the opposite of the top view. Rendering Lidar pointclouds from below allows to clearly see the colors of the flat surfaces of the ground without occlusions from the objects above.
We use all 4 types of camera poses, with 250 camera poses per type, per scene, thus 1,000 total views per scene. Using more views seems not to improve performance, and using fewer views leads to a drop.
III-D Multi-Modalities (MMods)
Similar to camera pose generation, PC2D datasets allow us to choose the visualization modality for the training and inference images. While RGB cameras are limited to RGB colors, we can choose to render modalities such as Lidar intensity, depth maps, and normals. The final projection and voting step III-E does not limit the number of votes, we therefore propose to train multiple models on multiple modalities, and use them together in an ensemble. We use the following modalities:
Intensities: The Lidar sensor intensity, normalized by scan.
RGB: AD Lidar datasets often provide RGB images, which can be used to colorize the pointclouds through projection. We render views of the obtained colored pointclouds.
Depths: The pointcloud depth maps with respect to the camera pose.
Normals: The components of the pointcloud surface normals, computed using the normal estimation method in [23].
For each modality, we generate a PC2D dataset, and train a separate model. We then use these 4 models to process each view of the scene at inference time, accumulating their segmentation outputs as votes. If a given dataset does not provide a modality (e.g. RGB), the remaining models can be used individually. The ablation study IV-F shows the performance is only slightly impacted when not using the RGB modality (-0.1 mIoU).
III-E Projection and Voting
We use a similar projection step as in [12], where 2D logits are back-projected to 3D and accumulated as votes. We propose adding occlusion to the projection step. Naive back-projection from 2D to 3D suffers from points being included in the projection even if they are occluded by other points. We use depth maps generated from meshes, themselves generated using [23] for occlusion. Using meshes has the advantage of allowing less sparse depth maps which is more appropriate for dense segmentation masks. Finally, we set a margin parameter to include points behind the depth map to account for the size of the occluding object.
IV Experiments
| NSSK10 (% IoU) | Car | Bicycle | Motorcycle | Truck | Other vehicle | Pedestrian | Driveable surf. | Sidewalk | Terrain | Vegetation | % mIoU | |
| Source-only | 74.09 | 0.01 | 10.18 | 1.07 | 1.0 | 4.43 | 63.44 | 31.81 | 36 | 25.6 | 24.76 | |
| T-UDA | [18] | 93.0 | 0.0 | 11.4 | 3.4 | 47.0 | 15.7 | 83.3 | 54.4 | 67.9 | 83.9 | 45.89 |
| Lidar-UDA | [19] | 86.19 | 0.0 | 13.87 | 9.30 | 3.1 | 16.49 | 65.69 | 6.07 | 54.05 | 85.65 | 37.48 |
| CoSMix | [20] | 77.1 | 10.4 | 20.0 | 15.2 | 6.6 | 51.0 | 52.1 | 31.8 | 34.5 | 84.8 | 38.35 |
| SALUDA | [21] | 89.8 | 13.2 | 26.2 | 15.3 | 7.0 | 37.6 | 79.0 | 50.4 | 55.0 | 88.3 | 46.18 |
| Seg_3D_by_PC2D | (ours) | 92.06 | 0.02 | 27.13 | 15.45 | 8.28 | 43.35 | 78.72 | 52.59 | 71.31 | 90.42 | 47.93 |
| Supervised | (SK10SK10) | 91.32 | 35.78 | 11.83 | 0.03 | 20.17 | 22.87 | 90.55 | 79.60 | 63.93 | 88.01 | 50.41 |
| SK10NS (% IoU) | Car | Bicycle | Motorcycle | Truck | Other vehicle | Pedestrian | Driveable surf. | Sidewalk | Terrain | Vegetation | % mIoU | |
| Source-only | 1.79 | 0 | 0 | 0.01 | 0 | 0.5 | 6.65 | 0.07 | 0.05 | 0.26 | 0.93 | |
| T-UDA | [18] | 74.2 | 0.5 | 40.3 | 21.8 | 0.2 | 0.4 | 87.8 | 45.8 | 46.1 | 70.3 | 43.02 |
| Lidar-UDA | [19] | 73.48 | 0.85 | 15.86 | 0.9 | 25.67 | 40.78 | 87.44 | 42.31 | 47.88 | 83.22 | 43.64 |
| Seg_3D_by_PC2D | (ours) | 72.2 | 0 | 44.67 | 49.47 | 0 | 27.79 | 90.03 | 56.06 | 59.57 | 87.88 | 48.77 |
| Supervised | (NSNS) | 71.96 | 20.16 | 29.87 | 36.52 | 0 | 55.20 | 93.28 | 72.11 | 65.73 | 86.14 | 58.38 |
In this section we demonstrate the performance and capabilities of our method on Lidar Semantic Segmentation tasks. In particular, the 2D reliance of our method shows the most benefits when applied to domain adaptation. To that end, the multi-view projection pipeline is used to generate pseudo-labels for the target dataset, which can be used to fine-tune a pre-trained 3D model. This indirect approach ensures a fair comparison by using similar model architectures as the state-of-the-art, and allows for faster downstream inference. In addition, the results are improved compared to the direct output of the pipeline, for reasons discussed in section IV-D. First, we show the performance of our method in the Real-to-Real setting, where we compare it to the state-of-the-art methods in unsupervised domain adaptation. Then, we apply our method to the Simulation-to-Real setting, where we show that our approach can reach state-of-the-art performance on large, structured classes, although not reaching the state-of-the-art when averaged over all classes. Then, we demonstrate how our method can be used to segment rare classes, for which 3D annotations are not available, by only using 2D annotations for those classes and 3D annotations for other classes. Finally, we conduct an ablation study to demonstrate the importance of each component of our method.
IV-A Experimental setup
IV-A1 Datasets
Real-to-Real
SemanticKITTI [3] (SK) and nuScenes [2] (NS) are two well-known large-scale datasets for AD, both providing and annotating Lidar 3D pointclouds. The procedure generally adopted in the domain adaptation literature [24, 25, 18, 19]
is to train on the official training splits for both datasets, e.g. 700 scenes for nuScenes and sequences 00-07 and 09-10 for SemanticKITTI, and evaluate on the official validation splits, e.g. 150 scenes for nuScenes and sequences 08 for SemanticKITTI. We follow this procedure for comparing with the state of the art. Of the 700 nuScenes scenes, we use the 70 last scenes for validation. Similarly, we use sequence 10 of SemanticKITTI for validation. For the rare classes experiment IV-E2 and the ablation study IV-F, we use a subset of the first 1800 scans from sequence 08 of the SK dataset for evaluation, to reduce the computational cost.
Simulation-to-Real
SynLidar [4] (SL) is a synthetic AD dataset captured with Unreal Engine 4. The sub-sampled version that is used in most of the literature [4, 20, 21, 22] is composed of 13 sequences, containing 19,865 scans with annotations for 32 classes. No official validation split is provided, therefore we follow [24] and use sequences 05 and 10 for validation. We do not evaluate on SynLidar, as it is only used as a source dataset.
IV-A2 Metrics
We measure the 3D semantic segmentation performance using the Intersection over Union (IoU) metric for all the experiments. We report both the per-class IoU and the mean IoU over all classes (mIoU). We compute all the metrics using code provided by [26], to ensure a fair comparison. For all UDA experiments, the evaluation is done on the 3D model after training on the pseudo-labels provided by Seg_3D_by_PC2D . For the rare classes and ablation experiments, the evaluation is directly done on the labels generated by the pipeline.
IV-B Implementation details
There are many hyper-parameters in our method that allow to try many different approaches. We present the most important ones here, and how we set them for best performance. Detailed parameters used will be available on github111https://github.com/andrewcaunes/ia4markings.
To begin with, we prepare the aligned scenes from SemanticKITTI and nuScenes. We use the already existing scenes from NS and accumulate scans 60 by 60 to create SK scenes. The intensities are normalized by scan, following [12]. The RGB images are used to colorize the pointclouds in both datasets using 2D-to-3D projection and a depth buffer to apply occlusion. Note that SynLidar does not provide RGB images, we therefore do not use RGB colors for the Sim-to-Real setting, where we use only 3 modalities: intensities, depths and normals.
NuScenes’ annotations are only available at 2Hz, while the scans are captured at 20Hz. To prepare for 2D segmentation label generation for PC2D datasets, we make the annotations denser by using a simple KNN method with to assign labels to each point in the dense scene.
The PC2D datasets are generated with 400,000 samples for each modality and source dataset. The views are uniformly sampled over the training scenes. For inference, as mentioned in III-C, we use 250 images per camera pose type, thus 1,000 images per scene, per modality. This amounts to 4,000 images total per scene. The inference time for each scene is relatively long, ~ 2 minutes on a NVIDIA RTX 3070 GPU, thus why the pseudo-label generation and fine-tuning application would be favored for real-time applications.
All views are rendered at 1024512 resolution with a standard field of view..
For the 2D segmentation models, we use the Mask2Former [27] model with Swin backbone [28], from the framework of [29]. We pretrain the model on the PC2D dataset for 70,000 iterations on 1024512 resolution images with a batch size of 8. We use a fading learning rate of .
For a fair comparison in UDA, we follow [18, 19, 21] and use as 3D model a MinkowskiNet 32 from [6] as the 3D segmentation model with the framework of [26]. We train it with the default settings from [26] for epochs on the source dataset, followed by iterations with batch size 4 on the target pseudo-annotated dataset. The only augmentations used are basic random 3D flip, rotation, translation and scaling.
IV-B1 Class mappings
We evaluate two Real-to-Real settings, NSSK and SKNS, and one Sim-to-Real setting, SLSK. To compare with the state-of-the-art, we follow [19, 21] and use 10 classes common to NS and SK. For the SLSK scenario, we use the official SK19 class mapping as in [20, 21]. To denote the different versions of the SemanticKITTI dataset, we use the following notations: SK10, SK19. The exact class mappings used can be found on github††footnotemark: .
| SLSK19 (% IoU) | Car | Bicycle | Motorcycle | Truck | Other vehicle | Pedestrian | Bicyclist | Motorcyclist | Road | Parking | Sidewalk | Other ground | Building | Fence | Vegetation | Trunk | Terrain | Pole | Traffic sign | % mIoU | |
| Source-only | 50.8 | 7.6 | 11.8 | 1.4 | 3.2 | 15.8 | 34.8 | 3.9 | 30.0 | 4.5 | 34.5 | 0 | 27.0 | 15.8 | 62.9 | 24.4 | 43.0 | 18.9 | 6.4 | 20.9 | |
| CoSMix | [20] | 75.1 | 6.8 | 29.4 | 27.1 | 11.1 | 22.1 | 25.0 | 24.7 | 79.3 | 14.9 | 46.7 | 0.1 | 53.4 | 13 | 67.7 | 31.4 | 32.1 | 37.9 | 13.4 | 32.2 |
| SALUDA | [21] | 67.0 | 7.7 | 14.4 | 1.3 | 5.2 | 24.1 | 52.6 | 2.7 | 52.5 | 10.5 | 44.1 | 0.4 | 51.8 | 13.6 | 69.7 | 40.5 | 56.5 | 45.0 | 14.3 | 30.2 |
| Unimix | [22] | 80.3 | 6.1 | 32.5 | 29.2 | 10.6 | 23.7 | 30 | 24.5 | 62.2 | 15.9 | 46.5 | 0.1 | 60.9 | 15.8 | 70.7 | 34.9 | 41.2 | 38.6 | 17.1 | 33.7 |
| Seg_3D_by_PC2D | (ours) | 83.5 | 0.9 | 6 | 10.4 | 3.0 | 19.8 | 13.5 | 0.2 | 80.2 | 0.0 | 51.6 | 0 | 78.8 | 16.7 | 51.9 | 32.0 | 36.8 | 40.6 | 10.1 | 28.2 |
IV-C Baselines
To demonstrate the capabilities of our method, we compare it to the state-of-the-art methods in UDA for 3D semantic segmentation. We use the results reported by the authors for each method, while ensuring a fair comparison by using the same dataset splits,class mappings, metrics and model architectures. For Real-to-Real, in the NSSK10 setting, we find the best performing methods to date are Lidar-UDA [19], T-UDA [18], CoSMix [20] and SALUDA [21]. In the SK10NS setting, we compare with Lidar-UDA [19] and T-UDA [18] as CoSMix [20] and SALUDA [21] do not provide results for this setting. Note that T-UDA [18] also reports values for the manmade class, not reported by the other methods. We therefore do not include it and re-compute the accordingly. For Sim-to-Real, in the SLSK19 setting, we compare with SALUDA [21], CoSMix [20] and UniMix [22].
IV-D Unsupervised Domain Adaptation Results
IV-D1 Real-to-Real setting
Results for the NSSK10 setting can be found in Table I. This UDA setting is particularly challenging as the source (NS) uses a 32-beam Lidar, less dense than the 64-beam Lidar used in the target (SK). Seg_3D_by_PC2D is able to outperform the state-of-the-art in mIoU by 1.75 Interestingly, our method reaches state-of-the-art performance even for dynamic classes such as car and motorcycle. This demonstrates that the model learns to match ‘shadow effect’-like features with the corresponding classes. Also reported in Table I is the supervised performance. This corresponds to the evaluation of the pseudo-labels outputed by the Seg_3D_by_PC2D pipeline using 2D models trained on the target dataset directly. It is interesting to note that in this ‘supervised’ setting, our method reaches relatively low performance compared to typical supervised methods (e.g. 65.3 mIoU in [21]), while still prevailing in UDA. We believe this is because the fully 2D reliance of the approach allows to bridge the domain adaptation gap successfully, while also hindering the potential performance by introducing errors in the projection and voting phase of the pipeline. In addition, it seems that the fine tuning of the pretrained model on pseudo-labels improves adaptation significantly, and the score of the pseudo-labels themselves does not act as a lower bound, indicating that there may exist a synergy between the source labels and target pseudo-labels that allows to reduce the domain gap.
Results for the SK10NS setting can be found in Table II. In this setting, Seg_3D_by_PC2D is even more dominant, with a 5.13 over state-of-the-art. In particular, the method is superior on all static classes, which can be explained by the higher difficulty to segment dynamic classes displaying some ‘shadow effect’. We notice relatively lower performance on small classes, such as bicycle and pedestrian, which intuitively is linked to the higher rate of precision errors in the projections.
IV-D2 Simulation-to-Real setting
Table III reports the results for the SLSK19 setting. Overall, our method achieves lower mIoU compared to prior works such as CoSMix [20], SALUDA [21], and UniMix [22]. However, it consistently performs strongly on several static classes, notably, road (80.24), sidewalk (51.55), and building (78.8), and some dynamic classes such as car (83.51), where it reaches state-of-the-art performance. This suggests that the projection-based pseudo-labeling is particularly effective on large, well-structured classes but less robust on fine-grained or rare classes such as bicycle, motorcycle, and motorcyclist. While the global mIoU is lower (28.21), the results highlight that our approach can still surpass prior methods on key static categories, underlining its strength for Sim-to-Real adaptation in structured environments
IV-E Application to rare classes
A very common real world scenario in 3D semantic segmentation is to have access to 3D annotations for some classes in a source domain, but not the desired target classes. However, 2D annotations are more generally available as they are less costly. We show that our method can be used in such cases: by leveraging the 3D annotations and our PC2D dataset generation pipeline, a 2D model is pretrained in the source domain. Then, using the 2D annotations, the model is fine-tuned on the target classes. This is an improvement over [12], which only uses 2D annotations. The pretraining on 3D annotations allows to significantly bridge the domain gap that exists between camera images and rendered images.
IV-E1 Experimental setup
For this experiment, we need 3 datasets:
A pretraining 3D source dataset, with 3D annotations for any classes. We use nuScenes.
A rare classes 2D dataset, with 2D annotations for the desired target classes. We use Mapillary Vistas (MV) [17], using only the traffic sign class as the target class.
A target 3D dataset. We use SemanticKITTI, as it conveniently provides 3D annotations for our target class, traffic sign, which will allow to perform a quantitative evaluation.
We compare Seg_3D_by_PC2D with seg_3D_by_2D from [12], trained on the Mapillary Vistas dataset. Differently from [12], we use the RGB colorized pointclouds for SK and NS instead of Lidar intensities, to reduce the domain gap with the MV dataset. We use Mask2Former models for both methods with the same parameters as IV-B. For Seg_3D_by_PC2D , we pretrain the 2D model on a PC2D dataset as in IV-B, and then fine-tune the model on the MV dataset for 10,000 iterations, while freezing the first block of the Swin backbone.
IV-E2 Results
Table IV reports the results as computed on SK for the traffic sign class. We see a significant improvement over the base method, with a 6.45 mIoU improvement. This shows that our method can be used to segment rare classes for which 3D annotations are not available.
| Method | traffic sign | ||
| Seg_3D_by_2D | [12] | 11.85 | |
| Seg_3D_by_PC2D | (ours) | 18.3 | +6.45 |
IV-F Ablation Study
We conduct a complete ablation study to verify the importance of each component of our method. Results can be seen in Table V. The first stage (a) corresponds to seg_3D_by_2D from [12]. We add the PC2D dataset training is stage (b), the occlusion handling in stage (c), the View Generation Optimization in stage (d), the Multi-Modalities module in stage (e), and the full Seg_3D_by_PC2D is evaluated in stage (f). We evaluate the models in both SK10SK10 (in-domain) and NSSK10 (UDA).
We see that the addition of the PC2D dataset leads to the most significant improvement, followed by the occlusion handling and the multi-modalities module. Interestingly, the addition of the VGO module leads to an improvement only for the SK10SK10 setting, and not for UDA. This might be because using more diverse camera poses increases the risk of encountering a large domain shift within one camera pose type. It can also be noted that the addition of the RGB colorization only lead to a small improvement, which shows the method is applicable for datasets lacking RGB images. Overall, all components proposed in our method seems valuable for either supervised or UDA applications.
| PC2D | OCL | VGO | MMods | RGB | SK10 | NS | |
| (a) | 15 | 15 | |||||
| (b) | ✓ | 31.1 | 28.9 | ||||
| (c) | ✓ | ✓ | 42.0 | 34.4 | |||
| (d) | ✓ | ✓ | ✓ | 43.7 | 34.3 | ||
| (e) | ✓ | ✓ | ✓ | ✓ | 49.3 | 40.9 | |
| (f) | ✓ | ✓ | ✓ | ✓ | ✓ | 49.4 | 41.1 |
V Conclusion
We have presented a novel multi-view projection framework for 3D semantic segmentation that is especially well suited for UDA. After generating large-scale synthetic 2D datasets (PC2D) from aligned Lidar scenes to train an ensemble of 2D segmentation models, our method leverages occlusion-aware back-projection and voting to produce high-quality 3D labels without requiring any target features or annotations.
Extensive experiments demonstrate that our approach achieves state-of-the-art performance in UDA and can be used to segment ‘rare’ classes, with only 2D annotations for the class and 3D annotations for other classes are available.
While our approach thoroughly explored problem mentionned in Section II-B of what views to use for inference, and which models to infer with, interesting research could be done exploring problem of how to optimally use the numerous outputs from the ensemble of models on all views to assign final 3D labels. In particular, our simple logits accumulation and voting scheme does not use important information such as spatial proximity of the points or class size statistics. Information like this and other might be used to train a model at outputting the final 3D segmentation mask with more insight than a simple voting scheme.
We hope that our work will inspire further research into this direction.
Acknowledgments
This project was provided with AI computing and storage resources by GENCI at IDRIS thanks to the grant 2025- AD011012128R4 on the supercomputer Jean Zay’s V100 and H100 partitions.
Large language models were used for spelling correction.

### v3

@@
 3D semantic segmentation plays a pivotal role in autonomous driving and road infrastructure analysis, yet state-of-the-art 3D models are prone to severe domain shift when deployed across different datasets.
-In this paper, we propose a novel multi-view projection framework that excels in unsupervised domain adaptation (UDA).
-Our approach first aligns Lidar scans into coherent 3D scenes and renders them from multiple virtual camera poses to create large-scale synthetic 2D datasets (PC2D) in various modalities.
-We then use them to train an ensemble of 2D segmentation models in pointcloud view domain on each modality.
-During inference, the models process hundreds of views per scene; the resulting logits are back-projected to 3D with an occlusion-aware voting scheme to generate final point-wise labels.
-These labels can be used directly or to fine-tune a 3D segmentation model in the target domain.
-We evaluate our approach in both Real-to-Real and Simulation-to-Real UDA settings.
-We achieve state-of-the-art results in UDA on Real-to-Real.
-We also demonstrate how our method can be used to segment ‘rare classes’, for which target 3D annotations are not available, by only using 2D annotations for those classes and leveraging 3D annotations for other classes in a source domain.
-I Introduction
+In this paper, we propose an Unsupervised Domain Adaptation approach where a 3D segmentation model is trained on the target dataset using pseudo-labels generated by a novel multi-view projection framework.
+Our approach first aligns Lidar scans into coherent 3D scenes and renders them from multiple virtual camera poses to create large-scale synthetic 2D semantic segmentation datasets in various modalities.
+The generated datasets are used to train an ensemble of 2D segmentation models in point cloud view domain on each modality.
+During inference, the models process a large amount of views per scene; the resulting logits are back-projected to 3D with a depth-aware voting scheme to generate final point-wise labels.
+These labels are then used to fine-tune a 3D segmentation model in the target domain.
+We evaluate our approach Real-to-Real on the nuScenes and SemanticKITTI datasets.
+We also evaluate it Simulation-to-Real with the SynLidar dataset.
+Our contributions are a novel method that achieves state-of-the-art results in Real-to-Real Unsupervised Domain Adaptation, and we also demonstrate an application of our method to segment ‘rare classes’, for which target 3D annotations are not available, by only using 2D annotations for those classes and leveraging 3D annotations for other classes in a source domain.
+1 Introduction
 3D semantic segmentation of Lidar scenes is crucial for applications such as autonomous driving (AD) and road infrastructure analysis.
@@
 if the data was acquired in a different location or with a different sensor.
-The main attempt to mitigate this problem is Unsupervised Domain Adaptation (UDA).
-Most existing methods are tailored for AD, thus making concessions on accuracy to guarantee real-time performance on single or few Lidar scans.
-We propose a method that leverages aligned scenes and larger computation times to perform 3D semantic segmentation. In particular, it reaches state-of-the-art-level performance in UDA.
-Our method belongs to the multi-view projection family, where 3D scenes are analyzed via multiple 2D representations.
-We position virtual cameras around aligned Lidar scenes to generate rendered views from various perspectives.
-While similar methods use camera images for training and inference [1], we create large-scale synthetic datasets by rendering both the 3D scenes and their corresponding ground truth annotations.
-This enables a 2D segmentation model to be trained on data closely matching the inference domain, allowing to freely choose the optimal parameters for selecting and rendering the views.
-The extensive use of 2D models leads to interesting domain adaptation performance.
-We demonstrate the effectiveness of our method for unsupervised domain adaptation in Real-to-Real tasks on the nuScenes [2] and SemanticKITTI [3] datasets, and in a Simulation-to-Real task on the SynLidar [4] dataset.
+The main approach to mitigate this problem is Unsupervised Domain Adaptation (UDA).
+Most existing methods are tailored for AD, thus making concessions on accuracy to guarantee real-time performance on single or few Lidar scans [24, 23, 18, 30].
+We propose a method that leverages aligned scenes to generate high quality pseudo-labels for the target domain, which are used to fine-tune a final 3D segmentation model pre-trained on the source domain.
+Our approach reaches state-of-the-art performance in UDA while introducing no computational overhead to the final 3D segmentation model.
+Our pseudo-label generation process belongs to the multi-view projection family of methods, where 3D scenes are analyzed via multiple 2D representations.
+We position virtual cameras around aligned Lidar scenes to generate rendered views (virtual camera images) from various perspectives.
+While similar methods use real camera images for training and inference [8], we create large-scale synthetic 2D semantic segmentation datasets by rendering both the 3D scenes and their corresponding ground truth annotations.
+This eliminates the need for additional camera sensors and allows to freely choose the optimal parameters for selecting and rendering the views.
+We demonstrate the effectiveness of our method for UDA in Real-to-Real tasks on the nuScenes (NS) [2] and SemanticKITTI (SK) [1] datasets (NSSK and SKNS), and in a Simulation-to-Real task on the SynLidar (SL) [29] dataset (SLSK).
 We also demonstrate a potential application of our method to segment ‘rare classes’, where 3D annotations are not available for the desired target classes, but 2D annotations are available along with 3D annotations for other classes.
@@
 Our main contributions are the following:
-A novel modular multi-view projection framework with dataset generation capabilities that can be used to explore many approaches and parameters for 3D semantic segmentation
-A state-of-the-art method for unsupervised domain adaptation in 3D semantic segmentation
-We demonstrate an application of our method to segment ‘rare classes’ (defined above)
-II Related works
-II-A 3D semantic segmentation
+A method for Unsupervised Domain Adaptation in 3D semantic segmentation reaching state-of-the-art performance in
+nuScenes SemanticKITTI and SemanticKITTI nuScenes
+A novel multi-view projection framework for pseudo-label generation, with dataset generation capabilities
+A demonstration of an application of our method to segment ‘rare classes’ for which 3D annotations are not available
+2 Related works
+2.1 3D semantic segmentation
 3D semantic segmentation is the task of assigning a semantic label to each point in a 3D point cloud.
@@
 3D models directly operate on the point cloud using 3D sparse convolutions [5], while 2D models re-use 2D image segmentation methods by projecting the point cloud to a 2D plane [14].
-Importantly, 3D models have been shown to suffer from domain shift
-to a higher degree than 2D models [8].
+Importantly, LiDAR-based 3D semantic segmentation models are particularly sensitive to domain shift due to sensor-specific sampling and sparsity effects, often more so than 2D vision models [24, 13].
 2D approaches generally project the point cloud to a single 2D image with channels representing various features, e.g.
-height (in Bird’s Eye View), intensity, density, etc.
-Multi-view projection methods try to leverage 2D models even further by repeatedly using the same model on multiple views of the same 3D scene.
-II-B Multi-view projection for segmentation
+height, intensity, density, etc.
+Various projection approaches exist, including Bird’s Eye View (BEV) [14] and Range View (RV) [28].
+Multi-view projection methods try to leverage 2D models even further by repeatedly using the same model on multiple views of the 3D scene.
+2.2 Multi-view projection for segmentation
 The goal of multi-view projection methods is to extract as much knowledge from a 2D model as possible for 3D segmentation.
@@
 The problem then decomposes itself into two parts:
-1.
-Which views to use for inference and with which model ?
-2.
+P1.
+Which projections to use for the views and with which inference model?
+P2.
 How to merge the 2D segmentation masks to get unique 3D labels?
@@
 It is generally solved using either a voting scheme [12, 17, 3], where either the 2D masks or the probabilities/logits of the model are accumulated on points as votes, or with more complex methods such as Neural Networks [22, 16, 26].
-Our proposed method uses a simple voting scheme on accumulated logits from a single 2D model.
-In this paper, we focus on the first problem.
+Our proposed method uses a fast voting scheme on accumulated logits from a single 2D model.
+In this paper, we focus on problem P1.
 Initially, [25] proposed multi-view projection for indoor RGB-D scenes, where the views are simply the RGB-D images themselves.
-For outdoor Lidar scenes, the same reasoning led [1] to use aligned RGB camera images as views, for which 2D annotated data is available [17].
-However, this approach is limited as it requires using additional sensors, and the performance is tied to the quality of the sensor alignment.
+For outdoor Lidar scenes, the same reasoning led [8] to use RGB camera images as views, with cameras calibrated to the Lidar [21].
+However, this approach is limited as it requires using additional sensors, and the performance is tied to the quality of the sensor calibration.
 Mitigating this, [3] proposed to infer on rendered views of 3D scenes with an out-of-domain 2D model trained on camera images.
 This has the advantage of not requiring 3D labels or additional sensors, but comes with several limitations:
-The domain gap between camera images and views of Lidar pointclouds is substantial, which negatively impacts 2D model performance.
-The multi-view aspect of the method is limited because the view generation method must align with the camera images, e.g., be from a road user’s perspective.
+The domain gap between camera images and views of Lidar point clouds is significant, which negatively impacts 2D model performance.
+The multi-view capability of the method is limited, as virtual camera placement must remain close to physically plausible viewpoints, e.g., those of real road users.
 Because aligning Lidar scans is essential for obtaining dense 3D scenes, the method is limited to static classes, as dynamic classes often appear with a so-called ‘shadow effect’ in the aligned scene, which is out of the camera images’ domain.
-We propose to solve all of these limitations by leveraging 3D annotations.
-This makes our method more dependent on 3D annotated data, but allows to reach significantly better performance and to tackle dynamic classes.
+We propose to solve these three limitations by leveraging 3D annotations.
+This makes our method more dependent on 3D annotated data, but allows it to achieve significantly better performance and handle dynamic classes.
 Closest to our approach, [12] proposes to use 3D annotations to generate 2D synthetic datasets to train a 2D model for downstream multi-view projection for indoor RGB-D scenes.
 Our method distinguishes itself in its application to outdoor Lidar scenes, where much additional processing is required to obtain satisfying views, e.g.
-scan aligning and meshing, as well as in the choices for the feature rendering and view selection.
-We also propose an ensembling mechanism for handling the modality choice, and demonstrate the potential of such methods for unsupervised domain adaptation.
-II-C Unsupervised Domain Adaptation
-Unsupervised Domain Adaptation (UDA) aims to adapt a model trained on a labeled source domain to perform well on an unlabeled target domain , where no ground-truth labels are available.
+scan aligning, as well as in the choices for the feature rendering and view selection.
+We propose an ensembling mechanism for handling the modality choice.
+We also demonstrate the potential of such methods for pseudo-label generation for unsupervised domain adaptation.
+2.3 Unsupervised Domain Adaptation
+UDA aims to adapt a model trained on a labeled source domain to perform well on an unlabeled target domain , where no ground-truth labels are available.
 Unlike supervised adaptation, only the source labels and unlabeled target samples can be exploited, while the evaluation is performed on a disjoint set .
@@
 Our method differs by relying entirely on a multi-view 2D projection pipeline to generate pseudo-labels for the target domain.
-These pseudo-labels are then used to fine-tune a 3D segmentation model.
-This fundamentally different approach allows to reach state-of-the-art performance in UDA, especially on large, structured classes.
-III Method
-III-A Base method
-An overview of our method, called Seg_3D_by_PC2D , is shown in Figure 1. We choose to use [12] as a base 3D multi-view projection method, which we enhance and extend.
-We provide a brief description of that method below.
-Here, a scene refers to the concatenation of multiple Lidar scans aligned in a common reference frame.
-Note that this concatenation yields denser scenes, but may introduce so called ‘shadow effects’ in the scene, due to dynamic classes moving during capture.
+These pseudo-labels are used to fine-tune a 3D segmentation model pretrained on the source domain.
+This approach reaches state-of-the-art performance in UDA.
+3 Method
+An overview of our method, which we call Seg3DbyPC2D , is shown in Fig.
 1.
-Train a 2D model for 2D semantic segmentation on natural camera images.
+To achieve UDA in 3D semantic segmentation, we simply pretrain a 3D model on the source dataset, then fine-tune it on the target dataset using pseudo-labels.
+Our main contribution is therefore our pseudo-label generation pipeline, which we describe in detail in this section.
+In the following, we use the terms ‘scene’ to refer to a point cloud resulting from the accumulation of Lidar scans into a common frame, and ‘view’ to refer to a 2D image rendered from a 3D scene using a virtual camera.
+3.1 Pseudo-label Generation
+Our method can be decomposed in the source training phase and the target inference phase.
+In both phases, we obtain 3D scenes and sample virtual camera poses to render views of the scenes.
+In the source training phase, we use these rendered views, along with rendered views of the 3D segmentation masks, to generate synthetic segmentation datasets (PC2D) that will be used to train an ensemble of 2D models.
+In the target inference phase, the views are used for inference with the previously trained ensemble of models, yielding pseudo-labels for the target domain.
+The choices of parameters for Virtual Camera Placement (VCP) and view rendering from Multiple Modalities (MMods) are crucial and are detailed in the following subsections.
+To obtain final pseudo-labels in the target domain, our multi-view projection approach is based on [3], which we improve with occlusion handling.
+3.1.1 Training with 3D annotations in 2D (PC2D)
+A major contribution to our work compared to existing multi-view projection methods is the training of 2D models in rendered view domain.
+To that end, we generate large-scale synthetic datasets of 2D views of 3D scenes along with 2D segmentation masks, as shown in Fig.
 2.
-Render views of a 3D scenes, then process them with the model, which is out of domain.
-3.
-Project the 2D segmentations on the 3D pointcloud and accumulate the logits, as votes.
-4.
-For each point, assign the most voted label.
-This method yields a 3D segmentation mask while using no 3D annotation at all.
-Our key innovation is to bridge the domain gap between camera images and rendered views of Lidar pointclouds by re-training the 2D model in pointcloud view domain, using 3D annotations.
-This allows to drastically improve the performance of the base method, and can still be applied without 3D annotations for the target classes, see the rare classes experiment IV-E2.
-First, we use 3D data along with 3D annotations to generate large-scale synthetic datasets of views of 3D scenes along with 2D segmentation masks, as shown in Figure 2, which we use to train the 2D model.
-We call these datasets ‘PointCloud2D’ datasets (PC2D).
-The parameters of the view generation are chosen carefully, as they are the same that will be used for inference downstream.
-We optimize the number and type of virtual camera poses (VGO, III-C), as well as the visualization modalities used.
-We use multiple models trained on each modality, and use them together in as an ensemble (MMods, III-D).
-Additionally, we improve the base method’s projection by adding occlusion using depth maps (OCL, III-E).
-In the experiments, we validate all of these components in the ablation study IV-F.
-III-B Training in pointcloud view domain using 3D annotations in 2D (PC2D)
-In [12], there is a substantial gap between the domain of the generated views and the domain of the camera training images.
-The main extension we propose to the base method is to train the 2D model in a much closer domain using 3D annotations.
-To this end, we generate a large-scale synthetic dataset of 2D views of 3D scenes along with 2D segmentation masks, as shown in Figure 2.
-Virtual camera poses are sampled around the 3D scene, which are then used to render 2D views of the scene.
-This way, we can generate an arbitrary number of (image, label) pairs, yielding a large-scale synthetic dataset.
-Training on such a synthetic dataset can be seen as a powerful data augmentation.
-The only limitation is that a limited number of annotated scenes will lead to samples with less diversity as the total number of training samples increases, which could lead to overfitting.
-The use of PC2D datasets allows to choose the domain of the training images, which allows full freedom over the view generation steps during both training and inference e.g.
-generating bird’s eye view depth maps is possible, while views in [12], were limited to ‘car’ view rendering of Lidar intensity.
-The two following subsections describe how we optimize the camera pose generation and the modality choice to obtain the best performance.
-III-C View Generation Optimization (VGO)
-We propose to optimize the view generation by choosing relevant camera poses.
-In [12], the camera poses are all placed at a road user’s perspective to match training data.
-We keep this category and add 3 new categories of camera poses.
-All 4 types of camera poses are obtained based on the original Lidar sensor poses, which are already describing the trajectory of the ego vehicle along the scene.
-An illustration of the 4 types of camera pose is shown in Figure 2, in the VGO module, and we describe them below:
-car_view: The camera is placed at a road user’s perspective, i.e. at the approximate height of a car (random yaw can be applied).
-conic_view: The camera is placed on the basis of a large imaginary cone pointing downwards, with the apex at the Lidar sensor position.
+We call these PC2D datasets.
+Virtual camera poses are sampled arbitrarily around the 3D scene, which are then used to render 2D views.
+This way, we can generate an arbitrary number of (image, label) pairs, yielding large-scale datasets.
+Training on such datasets can be seen as a powerful data augmentation technique.
+The main limitation to this is that a limited number of annotated scenes will yield samples with less diversity as the total number of training samples increases.
+The use of PC2D datasets allows to train the models in a rendered view source domain, very close to the target inference domain, yielding greater freedom over the view generation steps compared to existing methods.
+For example, generating bird’s eye view depth maps is possible, while views in [3], are limited to ‘road user viewpoint’ rendering of Lidar intensity.
+The two following subsections describe how we choose the virtual camera placement and the rendering modalities to obtain the best performance.
+3.1.2 Virtual Camera Placement
+We propose to optimize the view generation by choosing relevant virtual camera poses.
+After experimentation, we retain 4 types of camera poses based on the original Lidar sensor poses, which are already describing the trajectory of the ego vehicle along the scene.
+An illustration of the 4 types of camera pose is shown in Fig.
+2, in the VCP module, and we describe them below:
+carView: The camera is placed at a road user’s perspective, i.e. at the approximate height of a car (random yaw can be applied).
+conicView: The camera is placed on the basis of a large imaginary cone pointing downwards, with the apex at the Lidar sensor position.
 The camera points to the apex of the cone.
 This corresponds to a ‘drone’ or ‘lamp post’ view of the scene, and is also the closest to the views that human annotators would typically use to annotate a scene.
-top_view: The camera is placed directly above the Lidar position, looking downwards, with variable height and roll. This corresponds to the usual ‘bird’s eye view’ of the scene.
-bottom_view: The camera is placed directly below the Lidar position, looking upwards.
-This is the opposite of the top view.
+topView: The camera is placed directly above the Lidar position, looking downwards, with variable height and roll. This corresponds to the usual ‘bird’s eye view’ (BEV) of the scene.
+bottomView: The camera is placed under the scene, looking upwards.
+This is the opposite of a BEV.
 Rendering Lidar point clouds from below allows to clearly see the colors of the flat surfaces of the ground without occlusions from the objects above.
-We use all 4 types of camera poses, with 250 camera poses per type, per scene, thus 1,000 total views per scene.
-Using more views seems not to improve performance, and using fewer views leads to a drop.
-III-D Multi-Modalities (MMods)
-Similar to camera pose generation, PC2D datasets allow us to choose the visualization modality for the training and inference images.
-While RGB cameras are limited to RGB colors, we can choose to render modalities such as Lidar intensity, depth maps, and normals.
-The final projection and voting step III-E does not limit the number of votes, we therefore propose to train multiple models on multiple modalities, and use them together in an ensemble.
+3.1.3 Multi-Modalities
+Similar to camera pose generation, PC2D datasets allow us to choose the visualization modality for the training and inference images, i.e.
+the rendered colors.
+While multi-view projections methods based on RGB cameras are limited to RGB colors, we can choose to render modalities such as Lidar intensity, depth maps, and normals.
+The final projection and voting step does not limit the number of votes, we therefore propose to train multiple models on multiple modalities, and use them together in an ensemble.
 We use the following modalities:
-Intensities: The Lidar sensor intensity, normalized by scan.
-RGB: AD Lidar datasets often provide RGB images, which can be used to colorize the pointclouds through projection. We render views of the obtained colored pointclouds.
-Depths: The pointcloud depth maps with respect to the camera pose.
-Normals: The components of the pointcloud surface normals, computed using the normal estimation method in [23].
-For each modality, we generate a PC2D dataset, and train a separate model.
-We then use these 4 models to process each view of the scene at inference time, accumulating their segmentation outputs as votes.
-If a given dataset does not provide a modality (e.g.
-RGB), the remaining models can be used individually.
-The ablation study IV-F shows the performance is only slightly impacted when not using the RGB modality (-0.1 mIoU).
-III-E Projection and Voting
-We use a similar projection step as in [12], where 2D logits are back-projected to 3D and accumulated as votes.
-We propose adding occlusion to the projection step.
-Naive back-projection from 2D to 3D suffers from points being included in the projection even if they are occluded by other points.
+For each modality, we generate a PC2D dataset, and train a separate model. We then use ensembles of models to process each view of the scene at inference time.
+3.1.4 Projection and Fusion
+To obtain final 3D segmentation masks from 2D models’ outputs, we use a similar projection step as in [3], where 2D logits are back-projected to 3D and accumulated as votes.
+We propose adding occlusion to the projection step (OCL).
+Naive back-projection from 2D to 3D suffers from points being included in the projection even if they are occluded by objects.
 We use depth maps generated from meshes, themselves generated using [9] for occlusion.
-Using meshes has the advantage of allowing less sparse depth maps which is more appropriate for dense segmentation masks.
+Using meshes has the advantage of allowing less sparse depth maps which is more appropriate for dense point clouds.
 Finally, we set a margin parameter to include points behind the depth map to account for the size of the occluding object.
-IV Experiments
-| NSSK10 (% IoU) | Car | Bicycle | Motorcycle | Truck | Other vehicle | Pedestrian | Driveable surf. | Sidewalk | Terrain | Vegetation | % mIoU | |
-| Source-only | 74.09 | 0.01 | 10.18 | 1.07 | 1.0 | 4.43 | 63.44 | 31.81 | 36 | 25.6 | 24.76 | |
-| T-UDA | [18] | 93.0 | 0.0 | 11.4 | 3.4 | 47.0 | 15.7 | 83.3 | 54.4 | 67.9 | 83.9 | 45.89 |
-| Lidar-UDA | [19] | 86.19 | 0.0 | 13.87 | 9.30 | 3.1 | 16.49 | 65.69 | 6.07 | 54.05 | 85.65 | 37.48 |
-| CoSMix | [20] | 77.1 | 10.4 | 20.0 | 15.2 | 6.6 | 51.0 | 52.1 | 31.8 | 34.5 | 84.8 | 38.35 |
-| SALUDA | [21] | 89.8 | 13.2 | 26.2 | 15.3 | 7.0 | 37.6 | 79.0 | 50.4 | 55.0 | 88.3 | 46.18 |
-| Seg_3D_by_PC2D | (ours) | 92.06 | 0.02 | 27.13 | 15.45 | 8.28 | 43.35 | 78.72 | 52.59 | 71.31 | 90.42 | 47.93 |
-| Supervised | (SK10SK10) | 91.32 | 35.78 | 11.83 | 0.03 | 20.17 | 22.87 | 90.55 | 79.60 | 63.93 | 88.01 | 50.41 |
-| SK10NS (% IoU) | Car | Bicycle | Motorcycle | Truck | Other vehicle | Pedestrian | Driveable surf. | Sidewalk | Terrain | Vegetation | % mIoU | |
-| Source-only | 1.79 | 0 | 0 | 0.01 | 0 | 0.5 | 6.65 | 0.07 | 0.05 | 0.26 | 0.93 | |
-| T-UDA | [18] | 74.2 | 0.5 | 40.3 | 21.8 | 0.2 | 0.4 | 87.8 | 45.8 | 46.1 | 70.3 | 43.02 |
-| Lidar-UDA | [19] | 73.48 | 0.85 | 15.86 | 0.9 | 25.67 | 40.78 | 87.44 | 42.31 | 47.88 | 83.22 | 43.64 |
-| Seg_3D_by_PC2D | (ours) | 72.2 | 0 | 44.67 | 49.47 | 0 | 27.79 | 90.03 | 56.06 | 59.57 | 87.88 | 48.77 |
-| Supervised | (NSNS) | 71.96 | 20.16 | 29.87 | 36.52 | 0 | 55.20 | 93.28 | 72.11 | 65.73 | 86.14 | 58.38 |
-In this section we demonstrate the performance and capabilities of our method on Lidar Semantic Segmentation tasks.
-In particular, the 2D reliance of our method shows the most benefits when applied to domain adaptation.
-To that end, the multi-view projection pipeline is used to generate pseudo-labels for the target dataset, which can be used to fine-tune a pre-trained 3D model.
-This indirect approach ensures a fair comparison by using similar model architectures as the state-of-the-art, and allows for faster downstream inference.
-In addition, the results are improved compared to the direct output of the pipeline, for reasons discussed in section IV-D.
-First, we show the performance of our method in the Real-to-Real setting, where we compare it to the state-of-the-art methods in unsupervised domain adaptation.
-Then, we apply our method to the Simulation-to-Real setting, where we show that our approach can reach state-of-the-art performance on large, structured classes, although not reaching the state-of-the-art when averaged over all classes.
+This final step yields the pseudo-labels for the target dataset.
+4 Experiments
+In this section we demonstrate the performance of our method for UDA on Lidar Semantic Segmentation tasks.
+First, we show the performance of our method in both Real-to-Real and Simulation-to-Real settings.
+For both settings, we compare the results with the state-of-the-art methods in UDA.
 Then, we demonstrate how our method can be used to segment rare classes, for which 3D annotations are not available, by only using 2D annotations for those classes and 3D annotations for other classes.
 Finally, we conduct an ablation study to demonstrate the importance of each component of our method.
-IV-A Experimental setup
-IV-A1 Datasets
-Real-to-Real
+4.1 Datasets and Metrics
+Real-to-Real.
 SemanticKITTI [1] (SK) and nuScenes [2] (NS) are two well-known large-scale datasets for AD, both providing and annotating Lidar 3D point clouds.
@@
 is to train on the official training splits for both datasets, e.g.
-700 scenes for nuScenes and sequences 00-07 and 09-10 for SemanticKITTI, and evaluate on the official validation splits, e.g.
+700 scenes for nuScenes and sequences 00 to 07, 09 and 10 for SemanticKITTI, and evaluate on the official validation splits, e.g.
 150 scenes for nuScenes and sequences 08 for SemanticKITTI.
@@
 Similarly, we use sequence 10 of SemanticKITTI for validation.
-For the rare classes experiment IV-E2 and the ablation study IV-F, we use a subset of the first 1800 scans from sequence 08 of the SK dataset for evaluation, to reduce the computational cost.
-Simulation-to-Real
+For the rare classes experiment 4.6.2 and the ablation study 4.5, we use a subset of the first 1800 scans from sequence 08 of the SK dataset for evaluation, to reduce the computational cost.
+Simulation-to-Real.
 SynLidar [29] (SL) is a synthetic AD dataset captured with Unreal Engine 4.
@@
 We do not evaluate on SynLidar, as it is only used as a source dataset.
-IV-A2 Metrics
+Class mappings.
+We evaluate two Real-to-Real settings, NSSK and SKNS, and one Sim-to-Real setting, SLSK.
+To compare with the state-of-the-art, we follow [24, 18] and use 10 classes common to NS and SK.
+For the SLSK scenario, we use the official SK19 class mapping as in [23, 18].
+To denote the different versions of the SemanticKITTI dataset, we use the following notations: SK10, SK19.
+The exact class mappings used can be found on github111https://github.com/andrewcaunes/ia4markings.
+Metrics.
 We measure the 3D semantic segmentation performance using the Intersection over Union (IoU) metric for all the experiments.
@@
 We compute all the metrics using code provided by [19], to ensure a fair comparison.
-For all UDA experiments, the evaluation is done on the 3D model after training on the pseudo-labels provided by Seg_3D_by_PC2D .
-For the rare classes and ablation experiments, the evaluation is directly done on the labels generated by the pipeline.
-IV-B Implementation details
-There are many hyper-parameters in our method that allow to try many different approaches.
-We present the most important ones here, and how we set them for best performance.
-Detailed parameters used will be available on github111https://github.com/andrewcaunes/ia4markings.
+For all UDA experiments, the evaluation is done with the 3D model after fine-tuning, to ensure similar conditions to the state-of-the-art methods, i.e.
+real-time inference on single scans.
+For the rare classes and ablation experiments, the evaluation is directly done on the pseudo-labels generated by the pipeline.
+4.2 Implementation details
+The most important hyper-parameters are presented in this section, but more details can be found on github††footnotemark: .
 To begin with, we prepare the aligned scenes from SemanticKITTI and nuScenes.
@@
 The intensities are normalized by scan, following [3].
-The RGB images are used to colorize the pointclouds in both datasets using 2D-to-3D projection and a depth buffer to apply occlusion.
-Note that SynLidar does not provide RGB images, we therefore do not use RGB colors for the Sim-to-Real setting, where we use only 3 modalities: intensities, depths and normals.
 NuScenes’ annotations are only available at 2Hz, while the scans are captured at 20Hz.
-To prepare for 2D segmentation label generation for PC2D datasets, we make the annotations denser by using a simple KNN method with to assign labels to each point in the dense scene.
+To prepare for 2D segmentation label generation for PC2D datasets, we propagate the annotations to all points with a simple KNN method with to assign labels to each point in the dense scene.
 The PC2D datasets are generated with 400,000 samples for each modality and source dataset.
 The views are uniformly sampled over the training scenes.
-For inference, as mentioned in III-C, we use 250 images per camera pose type, thus 1,000 images per scene, per modality.
+For inference, we use 250 images per camera pose type, thus 1,000 images per scene, per modality.
 This amounts to 4,000 images total per scene.
-The inference time for each scene is relatively long, ~ 2 minutes on a NVIDIA RTX 3070 GPU, thus why the pseudo-label generation and fine-tuning application would be favored for real-time applications.
-All views are rendered at 1024512 resolution with a standard field of view..
+All views are rendered at 1024512 resolution with a standard field of view.
+The inference time for each scene takes ~ 2 minutes on a NVIDIA RTX 3070 GPU.
+This is however only done once for pseudo-label generation, and there is no overhead introduced to the final fine-tuned model in both memory and time.
+For the back-projection step, we crop the projection outside a depth range of [1m, 30m], and when occlusion is used, we set the occlusion margin to behind the depth encountered.
 For the 2D segmentation models, we use the Mask2Former [4] model with Swin backbone [15], from the framework of [20].
@@
 The only augmentations used are basic random 3D flip, rotation, translation and scaling.
-IV-B1 Class mappings
-We evaluate two Real-to-Real settings, NSSK and SKNS, and one Sim-to-Real setting, SLSK.
-To compare with the state-of-the-art, we follow [19, 21] and use 10 classes common to NS and SK.
-For the SLSK scenario, we use the official SK19 class mapping as in [20, 21].
-To denote the different versions of the SemanticKITTI dataset, we use the following notations: SK10, SK19.
-The exact class mappings used can be found on github††footnotemark: .
-| SLSK19 (% IoU) | Car | Bicycle | Motorcycle | Truck | Other vehicle | Pedestrian | Bicyclist | Motorcyclist | Road | Parking | Sidewalk | Other ground | Building | Fence | Vegetation | Trunk | Terrain | Pole | Traffic sign | % mIoU | |
-| Source-only | 50.8 | 7.6 | 11.8 | 1.4 | 3.2 | 15.8 | 34.8 | 3.9 | 30.0 | 4.5 | 34.5 | 0 | 27.0 | 15.8 | 62.9 | 24.4 | 43.0 | 18.9 | 6.4 | 20.9 | |
-| CoSMix | [20] | 75.1 | 6.8 | 29.4 | 27.1 | 11.1 | 22.1 | 25.0 | 24.7 | 79.3 | 14.9 | 46.7 | 0.1 | 53.4 | 13 | 67.7 | 31.4 | 32.1 | 37.9 | 13.4 | 32.2 |
-| SALUDA | [21] | 67.0 | 7.7 | 14.4 | 1.3 | 5.2 | 24.1 | 52.6 | 2.7 | 52.5 | 10.5 | 44.1 | 0.4 | 51.8 | 13.6 | 69.7 | 40.5 | 56.5 | 45.0 | 14.3 | 30.2 |
-| Unimix | [22] | 80.3 | 6.1 | 32.5 | 29.2 | 10.6 | 23.7 | 30 | 24.5 | 62.2 | 15.9 | 46.5 | 0.1 | 60.9 | 15.8 | 70.7 | 34.9 | 41.2 | 38.6 | 17.1 | 33.7 |
-| Seg_3D_by_PC2D | (ours) | 83.5 | 0.9 | 6 | 10.4 | 3.0 | 19.8 | 13.5 | 0.2 | 80.2 | 0.0 | 51.6 | 0 | 78.8 | 16.7 | 51.9 | 32.0 | 36.8 | 40.6 | 10.1 | 28.2 |
-IV-C Baselines
+4.3 Baselines
 To demonstrate the capabilities of our method, we compare it to the state-of-the-art methods in UDA for 3D semantic segmentation.
@@
 For Real-to-Real, in the NSSK10 setting, we find the best performing methods to date are Lidar-UDA [24], T-UDA [7], CoSMix [23] and SALUDA [18].
-In the SK10NS setting, we compare with Lidar-UDA [19] and T-UDA [18] as CoSMix [20] and SALUDA [21] do not provide results for this setting.
+In the SK10NS setting, we compare with Lidar-UDA [24] and T-UDA [7], as CoSMix [23] and SALUDA [18] do not provide results for this setting.
 Note that T-UDA [7] also reports values for the manmade class, not reported by the other methods.
-We therefore do not include it and re-compute the accordingly.
+We therefore do not include it and re-compute the mIoU accordingly.
 For Sim-to-Real, in the SLSK19 setting, we compare with SALUDA [18], CoSMix [23] and UniMix [30].
-IV-D Unsupervised Domain Adaptation Results
-IV-D1 Real-to-Real setting
-Results for the NSSK10 setting can be found in Table I.
+NSSK10 (% IoU) Car Bicycle Motorcycle Truck Other vehicle Pedestrian Driveable surf.
+Sidewalk Terrain Vegetation % mIoU Source-only 74.1 0.0 10.2 1.1 1.0 4.4 63.4 31.8 36.0 25.6 24.8 T-UDA [7] 93.0 0.0 11.4 3.4 47.0 15.7 83.3 54.4 67.9 83.9 46.0 Lidar-UDA [24] 86.2 0.0 13.9 9.3 3.1 16.5 65.7 6.1 54.1 85.7 34.1 CoSMix [23] 77.1 10.4 20.0 15.2 6.6 51.0 52.1 31.8 34.5 84.8 38.3 SALUDA [18] 89.8 13.2 26.2 15.3 7.0 37.6 79.0 50.4 55.0 88.3 46.2 Ours 90.1 0.0 28.4 17.7 18.4 37.1 80.9 61.0 70.2 90.1 49.4
+SK10NS (% IoU) Car Bicycle Motorcycle Truck Other vehicle Pedestrian Driveable surf.
+Sidewalk Terrain Vegetation % mIoU Source-only 1.8 0.0 0.0 0.0 0.0 0.5 6.7 0.1 0.1 0.3 0.9 T-UDA [7] 74.2 0.5 40.3 21.8 0.2 0.4 87.8 45.8 46.1 70.3 38.7 Lidar-UDA [24] 73.5 0.9 15.9 0.9 25.7 40.8 87.4 42.3 47.9 83.2 41.8 Ours 73.3 0.5 32.0 50.7 0.0 32.1 90.3 55.4 58.2 87.0 47.9
+4.4 Unsupervised Domain Adaptation Results
+4.4.1 Real-to-Real setting
+Results for the NSSK10 setting can be found in Table 1.
 This UDA setting is particularly challenging as the source (NS) uses a 32-beam Lidar, less dense than the 64-beam Lidar used in the target (SK).
-Seg_3D_by_PC2D is able to outperform the state-of-the-art in mIoU by 1.75 Interestingly, our method reaches state-of-the-art performance even for dynamic classes such as car and motorcycle.
+Seg3DbyPC2D is able to outperform the state-of-the-art by mIoU points.
+Interestingly, our method reaches state-of-the-art performance even for dynamic classes such as car, motorcycle and truck.
 This demonstrates that the model learns to match ‘shadow effect’-like features with the corresponding classes.
-Also reported in Table I is the supervised performance.
-This corresponds to the evaluation of the pseudo-labels outputed by the Seg_3D_by_PC2D pipeline using 2D models trained on the target dataset directly.
-It is interesting to note that in this ‘supervised’ setting, our method reaches relatively low performance compared to typical supervised methods (e.g.
-65.3 mIoU in [21]), while still prevailing in UDA.
-We believe this is because the fully 2D reliance of the approach allows to bridge the domain adaptation gap successfully, while also hindering the potential performance by introducing errors in the projection and voting phase of the pipeline.
-In addition, it seems that the fine tuning of the pretrained model on pseudo-labels improves adaptation significantly, and the score of the pseudo-labels themselves does not act as a lower bound, indicating that there may exist a synergy between the source labels and target pseudo-labels that allows to reduce the domain gap.
-Results for the SK10NS setting can be found in Table II.
-In this setting, Seg_3D_by_PC2D is even more dominant, with a 5.13 over state-of-the-art.
-In particular, the method is superior on all static classes, which can be explained by the higher difficulty to segment dynamic classes displaying some ‘shadow effect’.
-We notice relatively lower performance on small classes, such as bicycle and pedestrian, which intuitively is linked to the higher rate of precision errors in the projections.
-IV-D2 Simulation-to-Real setting
-Table III reports the results for the SLSK19 setting.
+Results for the SK10NS setting can be found in Table 2.
+In this setting, Seg3DbyPC2D is even more dominant, with a mIoU advantage over state-of-the-art.
+The method results in particularly strong performance on static classes such as driveable surface, sidewalk, terrain and vegetation, which can be explained by the higher difficulty to segment dynamic classes displaying some ‘shadow effect’.
+For both tasks, we notice relatively lower performance on small classes, such as bicycle and pedestrian, which might be due to the projection errors caused by poor sensor calibration being more important for small objects.
+SLSK19 (% IoU) Car Bicycle Motorcycle Truck Other vehicle Pedestrian Bicyclist Motorcyclist Road Parking Sidewalk Other ground Building Fence Vegetation Trunk Terrain Pole Traffic sign % mIoU Source-only 50.8 7.6 11.8 1.4 3.2 15.8 34.8 3.9 30.0 4.5 34.5 0 27.0 15.8 62.9 24.4 43.0 18.9 6.4 20.9 CoSMix [23] 75.1 6.8 29.4 27.1 11.1 22.1 25.0 24.7 79.3 14.9 46.7 0.1 53.4 13 67.7 31.4 32.1 37.9 13.4 32.2 SALUDA [18] 67.0 7.7 14.4 1.3 5.2 24.1 52.6 2.7 52.5 10.5 44.1 0.4 51.8 13.6 69.7 40.5 56.5 45.0 14.3 30.2 Unimix [30] 80.3 6.1 32.5 29.2 10.6 23.7 30 24.5 62.2 15.9 46.5 0.1 60.9 15.8 70.7 34.9 41.2 38.6 17.1 33.7 Ours 83.5 0.9 6 10.4 3.0 19.8 13.5 0.2 80.2 0.0 51.6 0 78.8 16.7 51.9 32.0 36.8 40.6 10.1 28.2
+4.4.2 Simulation-to-Real setting
+Table 3 reports the results for the SLSK19 setting.
 Overall, our method achieves lower mIoU compared to prior works such as CoSMix [23], SALUDA [18], and UniMix [30].
@@
 This suggests that the projection-based pseudo-labeling is particularly effective on large, well-structured classes but less robust on fine-grained or rare classes such as bicycle, motorcycle, and motorcyclist.
-While the global mIoU is lower (28.21), the results highlight that our approach can still surpass prior methods on key static categories, underlining its strength for Sim-to-Real adaptation in structured environments
-IV-E Application to rare classes
-A very common real world scenario in 3D semantic segmentation is to have access to 3D annotations for some classes in a source domain, but not the desired target classes.
+While the global mIoU is lower (28.2), the results highlight that our approach can still surpass prior methods on key static categories, underlining its strength for Sim-to-Real in some applications such as road infrastructure analysis.
+4.5 Ablation Study
+We conduct a complete ablation study to verify the importance of each component of our method.
+Results can be seen in Table 4.
+The first stage (a) corresponds to the base multi-view projection method that we followed from [3].
+We add the PC2D dataset training in stage (b), the occlusion handling in stage (c), the Virtual Camera Placement in stage (d), and the full Seg3DbyPC2D with the Multi-Modalities module is evaluated in stage (e).
+We evaluate the models in both SK10SK10 (in-domain) and NSSK10 (UDA).
+| PC2D | OCL | VCP | MMods | SK10 | NS | |||
+| (a) | 15 | - | 15 | - | ||||
+| (b) | ✓ | 31.1 | +16.1 | 28.9 | +13.9 | |||
+| (c) | ✓ | ✓ | 42.0 | +10.9 | 34.4 | +5.5 | ||
+| (d) | ✓ | ✓ | ✓ | 43.7 | +1.7 | 34.3 | -0.1 | |
+| (e) | ✓ | ✓ | ✓ | ✓ | 49.3 | +5.6 | 40.9 | +6.6 |
+We see that the addition of the PC2D dataset leads to the most significant improvement, followed by the occlusion handling and the multi-modalities module.
+Interestingly, the addition of the VCP module leads to an improvement only for the SK10SK10 setting, and not for UDA.
+This might be because using more diverse camera poses increases the risk of encountering a large domain shift within one camera pose type, e.g.
+the difference in Lidar point density might yield a larger difference in BEV type views than in road user type views, since the difference in point density becomes significant when seen from afar.
+Overall, all components proposed in our method seems valuable for either supervised or UDA applications.
+4.6 Application to rare classes
+A common real world scenario in 3D semantic segmentation is to have access to 3D annotations for some classes in a source domain, but not the desired target classes.
 However, 2D annotations are more generally available as they are less costly.
-We show that our method can be used in such cases: by leveraging the 3D annotations and our PC2D dataset generation pipeline, a 2D model is pretrained in the source domain.
-Then, using the 2D annotations, the model is fine-tuned on the target classes.
-This is an improvement over [12], which only uses 2D annotations.
-The pretraining on 3D annotations allows to significantly bridge the domain gap that exists between camera images and rendered images.
-IV-E1 Experimental setup
-For this experiment, we need 3 datasets:
-A pretraining 3D source dataset, with 3D annotations for any classes. We use nuScenes.
+For such a situation when 2D annotations of the target classes are available, [3] proposes a multi-view projection method with 2D models trained on real camera images, which then perform out-of-domain inference on virtual rendered views.
+We show that our method can help to significantly improve the performance of this approach by using the 3D annotations and our PC2D dataset generation pipeline to pretrain the 2D model in rendered view domain.
+The model is then fine-tuned on real camera images with annotations for the target classes.
+The pretraining on 3D annotations allows to significantly bridge the domain gap that exists at inference time by first exposing the model to rendered images.
+4.6.1 Experimental setup
+For this experiment, 3 datasets are used:
+A pretraining 3D source dataset, with 3D annotations for any classes (but not the target classes). We use nuScenes (NS) [2].
 A rare classes 2D dataset, with 2D annotations for the desired target classes. We use Mapillary Vistas (MV) [21], using only the traffic sign class as the target class.
-A target 3D dataset. We use SemanticKITTI, as it conveniently provides 3D annotations for our target class, traffic sign, which will allow to perform a quantitative evaluation.
-We compare Seg_3D_by_PC2D with seg_3D_by_2D from [12], trained on the Mapillary Vistas dataset.
-Differently from [12], we use the RGB colorized pointclouds for SK and NS instead of Lidar intensities, to reduce the domain gap with the MV dataset.
-We use Mask2Former models for both methods with the same parameters as IV-B.
-For Seg_3D_by_PC2D , we pretrain the 2D model on a PC2D dataset as in IV-B, and then fine-tune the model on the MV dataset for 10,000 iterations, while freezing the first block of the Swin backbone.
-IV-E2 Results
-Table IV reports the results as computed on SK for the traffic sign class.
-We see a significant improvement over the base method, with a 6.45 mIoU improvement.
-This shows that our method can be used to segment rare classes for which 3D annotations are not available.
+A target 3D dataset. We use SemanticKITTI (SK) [1], as it conveniently provides 3D annotations for our target class, traffic sign, which will allow to perform a quantitative evaluation.
+We compare our Seg3DbyPC2D with the baseline from [3], both trained on the rare classes dataset.
+Differently from [3] and from our base method in Sec.
+3, we use the RGB colorized point clouds using camera images for SK and NS as the rendered views’ modality instead of Lidar intensities or other modalities.
+This reduces the domain gap with the real camera images of the rare classes dataset, but requires that camera data be available for the pretraining dataset.
+We use Mask2Former models for both methods with the same parameters as 4.2.
+For Seg3DbyPC2D , we pretrain the 2D model on a PC2D dataset, generated from the pretraining dataset, as in 4.2, and then fine-tune the model on the rare classes dataset for 10,000 iterations, while freezing the first block of the Swin backbone.
 | Method | traffic sign | ||
-| Seg_3D_by_2D | [12] | 11.85 | |
-| Seg_3D_by_PC2D | (ours) | 18.3 | +6.45 |
-IV-F Ablation Study
-We conduct a complete ablation study to verify the importance of each component of our method.
-Results can be seen in Table V.
-The first stage (a) corresponds to seg_3D_by_2D from [12].
-We add the PC2D dataset training is stage (b), the occlusion handling in stage (c), the View Generation Optimization in stage (d), the Multi-Modalities module in stage (e), and the full Seg_3D_by_PC2D is evaluated in stage (f).
-We evaluate the models in both SK10SK10 (in-domain) and NSSK10 (UDA).
-We see that the addition of the PC2D dataset leads to the most significant improvement, followed by the occlusion handling and the multi-modalities module.
-Interestingly, the addition of the VGO module leads to an improvement only for the SK10SK10 setting, and not for UDA.
-This might be because using more diverse camera poses increases the risk of encountering a large domain shift within one camera pose type.
-It can also be noted that the addition of the RGB colorization only lead to a small improvement, which shows the method is applicable for datasets lacking RGB images.
-Overall, all components proposed in our method seems valuable for either supervised or UDA applications.
-| PC2D | OCL | VGO | MMods | RGB | SK10 | NS | |
-| (a) | 15 | 15 | |||||
-| (b) | ✓ | 31.1 | 28.9 | ||||
-| (c) | ✓ | ✓ | 42.0 | 34.4 | |||
-| (d) | ✓ | ✓ | ✓ | 43.7 | 34.3 | ||
-| (e) | ✓ | ✓ | ✓ | ✓ | 49.3 | 40.9 | |
-| (f) | ✓ | ✓ | ✓ | ✓ | ✓ | 49.4 | 41.1 |
-V Conclusion
-We have presented a novel multi-view projection framework for 3D semantic segmentation that is especially well suited for UDA.
-After generating large-scale synthetic 2D datasets (PC2D) from aligned Lidar scenes to train an ensemble of 2D segmentation models, our method leverages occlusion-aware back-projection and voting to produce high-quality 3D labels without requiring any target features or annotations.
-Extensive experiments demonstrate that our approach achieves state-of-the-art performance in UDA and can be used to segment ‘rare’ classes, with only 2D annotations for the class and 3D annotations for other classes are available.
-While our approach thoroughly explored problem mentionned in Section II-B of what views to use for inference, and which models to infer with, interesting research could be done exploring problem of how to optimally use the numerous outputs from the ensemble of models on all views to assign final 3D labels.
-In particular, our simple logits accumulation and voting scheme does not use important information such as spatial proximity of the points or class size statistics.
-Information like this and other might be used to train a model at outputting the final 3D segmentation mask with more insight than a simple voting scheme.
+| Baseline | [3] | 11.9 | |
+| Seg3DbyPC2D | (ours) | 18.3 | +6.4 (+53%) |
+4.6.2 Results
+Table 5 reports the results as computed on SK for the traffic sign class.
+We see a significant improvement over the base method, with a mIoU improvement, corresponding to relative increase of .
+This shows that our method can be used to segment rare classes for which 3D annotations are not available, when some 2D annotations are available.
+5 Conclusion
+We have presented a new UDA method based on a novel multi-view projection framework for 3D semantic segmentation.
+After generating large-scale synthetic 2D datasets (PC2D) from aligned Lidar scenes in the source domain to train an ensemble of 2D segmentation models, our method leverages occlusion-aware back-projection and voting to produce high-quality 3D pseudo-labels for the target domain.
+Extensive experiments demonstrate that our approach achieves state-of-the-art performance in Real-to-Real UDA and useful performance for large classes in Sim-to-Real UDA.
+We also demonstrated a potential application of our method to segment rare classes with no 3D annotations but 2D annotations available.
+While our approach thoroughly explored problem mentionned in Section 2.2 of what views to use for inference, and which models to infer with, future research could be led to explore problem of how to optimally use the numerous outputs from the ensemble of models on all views to assign final 3D labels.
+In particular, our logits accumulation and voting scheme does not use important information such as spatial proximity of the points or class size statistics.
+Information like this could be used to train a model to output the final 3D segmentation mask with more insight than a simple voting scheme.
 We hope that our work will inspire further research into this direction.
-Acknowledgments
-This project was provided with AI computing and storage resources by GENCI at IDRIS thanks to the grant 2025- AD011012128R4 on the supercomputer Jean Zay’s V100 and H100 partitions.
-Large language models were used for spelling correction.